Determine a hypothesis that the following data may address and perform a Chi-square test Goodness of Fit test on following data. (Outline H0 and H1, do the calculation--outline all your steps--and interpret the results, that is, link it back to either H0 or H1)

What is your political affiliation?

republican 102
democrat 98
total 200
Answer #1

Chi-Square Test

Null Hypothesis: Ho : There is no difference in the preference between republican versus democrat affiliation

Alternate hypothesis; Ha: There is significant difference in the preference between republican versus democrat affiliation

If the χ2 (calculated) < χ2 (table value) accept the null hypothesis at the 5% significant level otherwise reject the Ho and accept Ha.

Types of affiliation

Actual affiliation (O)

Expected Affiliation (E)

(O – E)2

(O – E)2/E

Republic

102

200/2 = 100

(102 – 100)2 = 4

4/100 = 0.04

Democrat

98

200/2 = 100

(98 – 100)2 = 4

4/100 = 0.04

χ2 = ∑[(O – E)2/E]

0.08

Degree of freedom: df = number of categories – 1 = 2 – 1 = 1

For significance of 0.05 and df = 1, χ2 (table value) = 3.841

As χ2 (calculated) < χ2 (table value), there is significant reason to accept the null hypothesis.

There is no difference in the preference between republican versus democrat affiliation